# A Lightweight Low-Light Image Enhancement Network via Channel Prior and Gamma Correction

### v1

@@ 4.2 Implementation Details
For LOLv2 and each fine-tuning stage of the training strategy for CPGA-Net (DGF), the networks were fine-tuned with 30 epochs and learning rate $10^{-5}$ using LOLv1 pretrained weights. CPGA-Net and CPGA-Net (DGF) employ different numbers of Resblock channels to improve performance and efficiency. In the regular version of CPGA-Net, we use 16 channels, while in CPGA-Net (DGF), we use 8 channels with a fast guided filter to reduce computational load and inference time.
@@ 4.4 Ablation study
Furthermore, our comprehensive analysis includes an extensive ablation study, as depicted in Fig. 8 and Table 5. The results of this study provide compelling evidence for the pivotal role of the perceptual loss in our model’s performance, it provides a substantial 0.16 dB improvement in PSNR and a 1% reduction in LPIPS, indicating that a significant enhancement in visual fidelity for both quantitative and perceptual aspects of image quality within our approach. For CPGA-Net (DGF), the results indicate that knowledge distillation significantly improves the overall performance, surpassing the performance without knowledge distillation and approaching the original performance achieved without utilizing the DGF module. This underscores the positive influence of knowledge distillation on the model’s performance and validates the interpretability of our model with knowledge distillation, paving the way for further extensions and applications.
@@
Table 5: Performance Analysis of Loss Functions on the LOLv1 dataset. Pre-tr. denotes as pretrained weight for fine-tuning, and the units of #Params. and FLOPs are M and G, respectively. (a) and (h) are the low-light image and normal-light image, respectively. (b) and (c) correspond to CPGA-Net with 16 channels of Resblocks, while (d) to (g) represent CPGA-Net with 8 channels of Resblocks. Specifically, (f) and (g) are instances of the CPGA-Net (DGF) structure with varying pretrained weights used for fine-tuning. The selected model of CPGA-Net and CPGA-Net (DGF) are (c) and (g), respectively. For each image quality metric, the top three data points of the experiments are shown in bold and underlined, bold only, and underlined only, respectively.
@@
(d) $\checkmark$ $\checkmark$ 20.31 0.70 0.28 0.018 4.36
(e) $\checkmark$ $\checkmark$ $\checkmark$ (d) 20.75 0.68 0.26 0.018 4.36
(f) $\checkmark$ $\checkmark$ (d) 19.79 0.70 0.30 0.018 1.10
(g) $\checkmark$ $\checkmark$ (e) 20.31 0.70 0.29 0.018 1.10
(h) $\infty$ 1.00 0.00 - -
Figure 7: Visualization results of the network architecture. (a) to (j) follow the order of Table 4, (k) represents the low-light image, and (l) represents the normal-light image.
Figure 8: Visualization results of loss functions. (a) to (h) follow the order of Table 5.

### v2

@@ 4.2 Implementation Details
For model initialization, we employ self-supervised training with 20 epochs, utilizing input images as targets to prime the model. This approach is exclusively applied to the train-from-scratch model. For LOLv2 and each fine-tuning stage of the training strategy for CPGA-Net (DGF), the networks were fine-tuned with 30 epochs and learning rate $10^{-5}$ using LOLv1 pretrained weights. CPGA-Net and CPGA-Net (DGF) employ different numbers of Resblock channels to improve performance and efficiency. In the regular version of CPGA-Net, we use 16 channels, while in CPGA-Net (DGF), we use 8 channels with a fast guided filter to reduce computational load and inference time.
@@ 4.4 Ablation study
Furthermore, our comprehensive analysis includes an extensive ablation study, as depicted in Fig. 8 and Table 5. The results of this study provide compelling evidence for the pivotal role of the perceptual loss in our model’s performance, it provides a substantial 0.16 dB improvement in PSNR and a 1% reduction in LPIPS, indicating that a significant enhancement in visual fidelity for both quantitative and perceptual aspects of image quality within our approach. In addition, self-supervised training improves performance with a 0.94 dB increase in PSNR, a 4% improvement in SSIM, and a 2% reduction in LPIPS. For CPGA-Net (DGF), the results indicate that knowledge distillation significantly improves overall performance, surpassing the performance without knowledge distillation and approaching the original performance achieved without utilizing the DGF module. This underscores the positive influence of knowledge distillation on the model’s performance and validates the interpretability of our model with knowledge distillation, paving the way for further extensions and applications. Both self-supervised training and knowledge distillation demonstrate the powerful capabilities of pretraining for deep learning image enhancement.
@@
Table 5: Performance Analysis of Loss Functions on the LOLv1 dataset. Pre-tr. denotes as pretrained weight for fine-tuning, and the units of #Params. and FLOPs are M and G, respectively. SST stands for self-supervised training. (a) and (i) are the low-light image and normal-light image, respectively. (b) to (d) correspond to CPGA-Net with 16 channels of Resblocks, while (e) to (h) represent CPGA-Net with 8 channels of Resblocks. Specifically, (g) and (h) are instances of the CPGA-Net (DGF) structure with varying pretrained weights used for fine-tuning. The selected model of CPGA-Net and CPGA-Net (DGF) are (d) and (h), respectively. For each image quality metric, the top three data points of the experiments are shown in bold and underlined, bold only, and underlined only, respectively.
@@
(d) $\checkmark$ $\checkmark$ SST 20.94 0.75 0.26 0.025 6.03
(e) $\checkmark$ $\checkmark$ 20.31 0.70 0.28 0.018 4.36
(f) $\checkmark$ $\checkmark$ $\checkmark$ (e) 20.75 0.68 0.26 0.018 4.36
(g) $\checkmark$ $\checkmark$ (e) 19.79 0.70 0.30 0.018 1.10
(h) $\checkmark$ $\checkmark$ (f) 20.31 0.70 0.29 0.018 1.10
(i) $\infty$ 1.00 0.00 - -
Figure 7: Visualization results of the network architecture ablation study. (a) to (j) follow the order of Table 4, (k) represents the low-light image, and (l) represents the normal-light image.
Figure 8: Visualization results of loss functions ablation study. (a) to (h) follow the order of Table 5.